Kim Fernando has just found the variance of a sample of 32 observations to be 64.0 by using a formula that is programmed into her new calculator. Therefore, what is the sample standard deviation?
Answer by RAY100(1637)

sample variance = s^2 = 64
.
sample standard deviation = s = sqrt(s^2) = sqrt(64) = 8
.